# On Singular Integrals and Quantitative Rectifiability in Parabolic Space and the Heisenberg Group

### v1

@@ 5.1. The parabolic case
For every $\lambda>0$ there exists $M_{2}>0$ such that the following holds: Suppose that $\mu$ is a $(k+2)$-ADR vampiric measure such that, there is a $k$-dimensional linear subspace $L\subset\operatorname{supp}(\mu)_{0}$ and vectors $e_{1},e_{2}$ orthogonal to $L$ with $|e_{1}|\leq 3\sqrt{n}$ with
(5.10) $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}\bigl(\mathbf{0},\frac{|e_{2}|}{\sqrt{n}}\bigl)\subset\operatorname{span}(L,e_{1})\cap\widetilde{Q}\bigl(\mathbf{0},\frac{|e_{2}|}{\sqrt{n}}\bigl),$
@@
(5.12) $\operatorname{supp}(\mu)_{0}\supset\Bigl\{L+\sum_{j=1}^{2}a_{j}e_{j}\text{ for any }a_{j}\in 4\mathbb{Z}\Bigl\}.$
@@
(5.13) $\alpha_{\mu,k+1}(B(\mathbf{0},M_{1}))<\lambda\text{ or }\alpha_{\mu,k+2}(B(\mathbf{0},M_{2}))<\lambda.$
@@
Suppose that the statement is false, so there exists $\lambda>0$ such that there is no $M_{2}(\lambda)$ for which the statement holds. Then for every (sufficiently large) $\ell$ there exists a $(k+2)$-ADR vampiric measure $\widetilde{\mu}_{\ell}$, a $k$-dimensional subspace $L_{\ell}$, and orthogonal vectors $e^{\ell}_{1}$ and $e^{\ell}_{2}$ with $|e_{1}^{\ell}|\leq 1$ and (5.10)–(5.12) hold, but (5.13) fails, with $M_{2}=\ell$. Recall $M_{1}=M_{1}(\lambda)$ from Lemma 5.12. If $|e_{2}^{\ell}|>5M_{1}\sqrt{n}$, then we observe from (5.12) that $\operatorname{supp}(\widetilde{\mu}_{\ell})_{0}\cap\widetilde{Q}(\mathbf{0},5M_{1})\subset\text{span}(L_{\ell},e_{1}^{\ell})\cap\widetilde{Q}(\mathbf{0},5M_{1})$, and so we can apply Lemma 5.12 to deduce that $\alpha_{\mu,k+1}(B(\mathbf{0},M_{1}))<\lambda$ (and therefore (5.13) holds), so we may assume that $|e_{2}^{\ell}|\leq 5M_{1}\sqrt{n}$.
Now we are in a position to repeat the analysis in Lemma 5.12. Consider the measures $\mu_{\ell}=\frac{\widetilde{\mu}_{\ell}(\,\cdot\,\ell)}{\ell^{k+2}}$. These measures are $(k+2)$-ADR and vampiric, so by passing to a subsequence we may assume converge weakly to an ADR measure (Lemma 2.1) $\mu$ that is vampiric. By weak continuity of $\alpha$-numbers (Lemma 2.7), we have that $\alpha_{\mu,k+2}(Q(\mathbf{0},1))\geq\lambda$. In $\widetilde{Q}(\mathbf{0},5)$, we obtain from (5.9) that $\operatorname{supp}(\mu_{\ell})_{0}$ contains $\bigl\{L+\frac{a_{1}}{\ell}e_{1}+\frac{a_{2}}{\ell}e_{2}:a_{1},a_{2}\in 4\mathbb{Z}\bigl\}\cap\widetilde{Q}(\mathbf{0},5)$. From this, along with the Ahlfors regularity of $\mu_{\ell}$, we conclude that $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}(\mathbf{0},5)\supset P\cap\widetilde{Q}(\mathbf{0},5)$, where $P=\text{span}(L,e_{1},e_{2})$, but additionally, from (5.11) we find that $\operatorname{supp}(\mu)\subset P$ so $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}(\mathbf{0},5)=P\cap\widetilde{Q}(\mathbf{0},5)$. Observe that $P$ is a $(k+2)$-plane. But now from Lemma 5.9, we have that, in $Q(\mathbf{0},1)$,
@@
but then $\alpha_{\mu,k+2}(B(\mathbf{0},1))=0$, which is absurd. ∎
We may now complete the
@@
The proof is a case analysis based upon Lemma 5.11.
Suppose $\mu$ is an $(n+2)$-ADR vampiric measure with $\alpha_{\mu,k}(B(\mathbf{0},1))>0.$. Consider the linear subspace $L\subset\operatorname{supp}(\mu)_{0}$ of dimension $m\in\{k,k+1,k+2\}$ given in Lemma 5.11.
First suppose that $\dim L=k$. Then there exists $e_{1}\perp L$, $|e_{1}|<3\sqrt{n}$ with $L+e_{1}\subset\operatorname{supp}(\mu)_{0}$. Indeed, otherwise $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}(\mathbf{0},3)=L$, and so Lemma 5.9 ensures that $\mu|_{Q(\mathbf{0},r)}=(\mathcal{H}^{k}|_{L}\times\nu)|_{Q(\mathbf{0},r)},$ which yields that $\alpha_{\mu,k}(B(\mathbf{0},1))=0$, a contradiction. Now, depending on whether or not there is a further orthogonal vector $e_{2}$, either we can apply Lemma 5.12 to get that $\alpha_{\mu,n+1}(B(\mathbf{0},M_{1}))<\lambda$, or Lemma 5.12 which yields (5.13), as required.
Next suppose $\dim L=k+1$. If $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}(\mathbf{0},5M_{1})=L$, then Lemma 5.9 ensures that $\mu|_{Q(\mathbf{0},M_{1})}=(\mathcal{H}^{k+1}|_{L}\times\nu)|_{Q(\mathbf{0},M_{1})},$ which yields that $\alpha_{\mu,k+1}(B(\mathbf{0},M_{1}))=0$. Otherwise, Lemma 5.11 ensures that there exists $e_{2}\in\operatorname{supp}(\mu)_{0}$ orthogonal to $L$. In this case we apply Lemma 5.13 (writing $L=\text{span}(\widetilde{L},e_{1})$ with $|e_{1}|<3\sqrt{n}$) to get that (5.13) holds. The conclusions have been verified again.
Finally, suppose $\dim L=k+2$. Then we must have $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}(\mathbf{0},5M_{2})=L$, and so Lemma 5.9 ensures that $\mu|_{Q(\mathbf{0},M_{2})}=(\mathcal{H}^{k+1}_{|L}\times\nu)|_{Q(\mathbf{X},M_{2})},$ which yields that $\alpha_{\mu,k+2}(B(\mathbf{0},M_{2}))=0$.
The proposition has been proved.∎
@@ 5.2. The Heisenberg group case
$\Sigma_{\mathbf{X}}(\mathbf{Y})=(2X-Y,t_{Y}+X_{2}Y_{1}-Y_{2}X_{1})$
@@
$=\mu(\mathbf{X}\cdot B_{r}(Y-X,t_{Y}-t_{X}+\frac{1}{2}(X_{2}Y_{1}-X_{1}Y_{2}))$
@@
Orponen [Orp] called a Borel measure $\mu$ on $\mathbb{H}$ symmetric measure if, for each pair $\mathbf{X},\mathbf{Y}\in\operatorname{supp}(\mu)$, $\Sigma_{\mathbf{X}}(\mathbf{Y})\in\operatorname{supp}(\mu)$, and proved the following result:
@@
[Orp] The support of a $3$-regular, symmetric measure on $\mu$ on $\mathbb{H}$ is contained in a vertical plane.
@@
Suppose $\mu$ is a $3$-regular, vampiric measure on $\mathbb{H}$ whose support contains $(0,0,0)$. Then $\mu=\mathcal{H}^{1}|_{L}\times\nu$ for some line $L$ and a $2$-ADR measure $\nu$ on $\mathbb{R}.$ Conversely, any measure of the form $\mu=\mathcal{H}^{1}|_{L}\times\nu$ is a vampiric measure on $\mathbb{H}$.
@@
Orponen’s theorem (Corollary 5.16 ensures that, for some vertical hyperplane $L\subset\mathbb{H}$, $\operatorname{supp}(\mu)\subset L$. If $\mathbf{X},\mathbf{Y}\in L$, then $\mathbf{X}\cdot\mathbf{Y}=(X+Y,t_{X}+t_{Y})$, and so the Heisenberg group action becomes the Parabolic group action on $L$. Therefore Lemma 5.11 ensures that $(\operatorname{supp}(\mu))_{0}=L$. Appealing now to Lemma 5.9 completes the proof. ∎
@@ 6. Packing bad cubes
Our main proposition is the following. Recall the quantities $M_{1}(\lambda)$ and $M_{2}(\lambda)$ in Proposition 5.2.
@@
(6.1) $\alpha_{\mu,k}(B(\mathbf{Z}_{E},\ell(E)))\geq\varkappa,\;\alpha_{\mu,k+1}(B(\mathbf{Z}_{E},M_{1}\ell(E))\geq\lambda,\,\text{ and }\alpha_{\mu,k+1}(B(\mathbf{Z}_{E},M_{2}\ell(E))\geq\lambda,$
@@
$\max_{f\in\mathcal{F}}\Theta_{\mu,\varphi,A}(E)\geq\Delta\mu(E).$
@@
Suppose that $\mu$ is $(k+2)$-ADR, and there exists $\lambda$ such that $\lambda$-$\bf HSDC$} cubes are Carleson. Then the collection of cubes $E\in\mathbb{D}(\mu)$ such that
$\alpha_{\mu,k+1}(B(\mathbf{Z}_{E},M_{1}\ell(E)))<\lambda\text{ or }\alpha_{\mu,k+2}(B(\mathbf{Z}_{E},M_{2}\ell(E)))<\lambda$
@@
Suppose that $\mu$ is a $(k+2)$-ADR measure for which is good for all spatially asymmetric Littlewood-Paley kernels, and the set of $\lambda$-$\bf HSDC$} cubes are Carleson for some $\lambda>0$. Fix $\varkappa$ to be equal to be the value of $\lambda$ in Theorem 4.3. First, using Lemma 2.5, we find that, for every the collection cubes $E\in\mathbb{D}(\mu)$ which satisfy (6.1) are Carleson. But now, since the $\lambda$-HSDC cubes are Carleson, Proposition 6.1 and Lemma 6.2 ensure that the collection of cubes $E\in\mathbb{D}(\mu)$ such that $\alpha_{\mu,k}(B(\mathbf{Z}_{E},\ell(E)))>\varkappa$ is Carleson. But now Theorem 4.3 ensures that $\mu$ is uniformly recitifiable.∎
@@
Fix $\varkappa>0$ and $\lambda>0$. Let $\{\varphi_{j}\}_{j}$ be a countable, dense subset of $\operatorname{Lip}_{0}^{\operatorname{odd}}(\mathbb{R}^{n+1})$ with $\|\varphi\|_{lip}\leq 1$, such that for each $j$, $\varphi_{j}\in C^{\infty}(\mathbb{R}^{n+1})$ and $\operatorname{supp}(\varphi_{j})\subseteq B_{j}(\mathbf{0}))$ for each $j$.
@@
We let $\mathbb{D}(\mu_{k})$ be the dyadic grid which is the image of $\mathbb{D}(\widetilde{\mu}_{k})$ under the mapping
@@
$\max_{j\in\{1,2,...,\ell\}}\Theta^{\mu_{\ell}}_{\varphi_{j},\ell}(Q_{\ell})\leq\frac{1}{\ell}$
@@
$\alpha_{\mu,k}(B(\mathbf{Z}_{E},\ell(E)))>\lambda,$

### v2

@@ 5.1. The parabolic case
For every $\lambda>0$ there exists $M_{2}>0$ such that the following holds for every $M\geq M_{2}$: Suppose that $\mu$ is a $(k+2)$-ADR vampiric measure such that $\mathbf{0}\in\operatorname{supp}(\mu)$ and there is a $k$-dimensional linear subspace $L\subset\operatorname{supp}(\mu)_{0}$ and linearly independent vectors $e_{1},e_{2}$ which are orthogonal to $L$ with $|e_{1}|\leq 3\sqrt{n}$ satisfying
(5.10) $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{|e_{2}|/\sqrt{n}}\bigl(0\bigl)\subset\operatorname{span}(L,e_{1})\cap\widetilde{Q}_{|e_{2}|/\sqrt{n}}\bigl(0\bigl),$
@@
(5.12) $\operatorname{supp}(\mu)_{0}\supset\Bigl\{L+\sum_{j=1}^{2}a_{j}e_{j}\text{ for any }a_{j}\in 4\mathbb{Z}\Bigl\},$
@@
(5.13) $\alpha_{k+1,\mu}(\mathbf{0},M_{1})<\lambda\text{ or }\alpha_{k+2,\mu}(\mathbf{0},M)<\lambda,$
where $M_{1}=M_{1}(\lambda)$ is the constant produced in Lemma 5.12.
@@
Suppose that the statement is false. Then there exists some $\lambda>0$ and a sequence of $(k+2)$-ADR vampiric measures measures $\{\widetilde{\mu}_{\ell}\}_{\ell\in\mathbb{Z}}$ with the following properties.
  • There exists an increasing sequence $\{M_{2,\ell}\}_{\ell\in\mathbb{N}}$ which tends to infinity such that, for each $l\in\mathbb{N}$,
    $\alpha_{k+1,\widetilde{\mu}_{\ell}}(\mathbf{0},M_{1}(\lambda))\geq\lambda\text{ and }\alpha_{k+2,\widetilde{\mu}_{\ell}}(\mathbf{0},M_{2,\ell})\geq\lambda.$
  • There exists a sequence of $k$-dimensional subspaces $\{L_{\ell}\}_{\ell\in\mathbb{N}}$ with $L_{\ell}\subset\operatorname{supp}(\widetilde{\mu}_{\ell})_{0}$ and sequences $\{e_{1,\ell}\}_{\ell\in\mathbb{N}}$, $\{e_{2,\ell}\}_{\ell\in\mathbb{N}}$, with $e_{i,\ell}\perp L_{\ell}$ for $i\in\{1,2\}$, and such that $|e_{1,\ell}|\leq 3\sqrt{n}.$
  • Displays 5.105.12 hold with $\mu$, $L$, $e_{1}$, and $e_{2}$ replaced by $\widetilde{\mu}_{\ell}$, $L_{\ell}$, $e_{1,\ell}$, $e_{2,\ell}$.
If $|e^{\ell}_{2}|>5M_{1}\sqrt{n}$ for infinitely many $l$, then, for such $\ell$, we observe from 5.10 that
$\operatorname{supp}(\widetilde{\mu}_{\ell})_{0}\cap\widetilde{Q}_{5M_{1}}(\mathbf{0})\subset\text{span}(L_{\ell},e_{1}^{\ell})\cap\widetilde{Q}_{5M_{1}}(\mathbf{0})$
and hence
$\widetilde{Q}_{5M_{1}}(\mathbf{0})\cap\{L_{l}+ae_{1}^{\ell},\,\,\forall a\in 4\mathbb{Z}\}\subset\widetilde{Q}_{5M_{1}}(\mathbf{0})\cap\operatorname{supp}(\mu_{\ell})_{0}.$
So, we can apply Lemma 5.12 to show that $\alpha_{k+1,\widetilde{\mu}_{\ell}}(0,M_{1})<\lambda.$ We are free to pass to a subsequence such that $|e^{\ell}_{2}|>5M_{1}\sqrt{n}$ for all $\ell$ and which is weakly convergent to a measure $\mu$ which, by Lemma 2.1. This measure will satisfy the hypotheses of the Lemma, and will (by Lemma 2.7) satisfy $\alpha_{k+1,\mu}(\mathbf{0},M_{1})<\lambda.$ Thus, we may assume that $|e_{2}^{\ell}|>5M_{1}\sqrt{n}$ for only finitely many $n$, and (after passing to a subsequence), assume that $|e_{2}^{\ell}|\leq 5M_{1}\sqrt{n}$ for all $n$.
Now we are in a position to repeat the analysis in Lemma 5.12. Consider the measures $\mu_{\ell}=\frac{\widetilde{\mu}_{\ell}(\,\cdot\,M_{2,\ell})}{M_{2,\ell}^{k+2}}$. These measures are $(k+2)$-ADR and vampiric, so by passing to a subsequence we may assume converge weakly to an ADR measure (Lemma 2.1) $\mu$ that is vampiric. We are also free to pass to a subsequence so that $e_{i,\ell}/|e_{i,\ell}|\rightarrow e_{i}$ for $i=1,2$ and so that $L_{l}$ converges to a $k$-plane $L$ in the local Hausdorff metric. By weak continuity of $\alpha$-numbers (Lemma 2.7), we have that $\alpha_{\mu,k+2}(Q_{1}(\mathbf{0}))\geq\lambda$. In $\widetilde{Q}_{5}(\mathbf{0})$, we obtain from (5.9) that $\operatorname{supp}(\mu_{\ell})_{0}$ contains $\bigl\{L+\frac{a_{1}}{M_{2,\ell}}e_{1}+\frac{a_{2}}{M_{2,\ell}}e_{2}:a_{1},a_{2}\in 4\mathbb{Z}\bigl\}\cap\widetilde{Q}_{5}(\mathbf{0})$. From this, along with the Ahlfors regularity of $\mu_{\ell}$, we conclude that $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{5}(\mathbf{0})\supset P\cap\widetilde{Q}_{5}(\mathbf{0})$, where $P=\text{span}(L,e_{1},e_{2})$, but additionally, from (5.11) we find that $\operatorname{supp}(\mu)\subset P$ so $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{5}(\mathbf{0})=P\cap\widetilde{Q}_{5}(\mathbf{0})$. Observe that $P$ is a $(k+2)$-dimensional plane. But now from Lemma 5.9, we have that, in $Q_{1}(\mathbf{0})$,
@@
but then $\alpha_{k+2,\mu}(\mathbf{0},1)=0$, which is absurd. ∎
We may now complete the proof of Proposition 5.2.
@@
The proof is a case analysis based upon Lemma 5.11. We rescale to assume, without loss of generality, that $r=1$.
Suppose $\mu$ is an $(k+2)$-ADR vampiric measure with $\alpha_{k,\mu}(\mathbf{0},1)>0$. Consider the linear subspace $L\subset\operatorname{supp}(\mu)_{0}$ of dimension $m\in\{k,k+1,k+2\}$ given in Lemma 5.11.
First suppose that $\dim L=k$. Then, there exists $e_{1}\perp L$, $|e_{1}|<3\sqrt{n}$ with $L+e_{1}\subset\operatorname{supp}(\mu)_{0}$. Indeed, otherwise $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{3}(\mathbf{0})=L$, and so Lemma 5.9 ensures that $\mu|_{Q_{3}(\mathbf{0})}=(\mathcal{H}^{k}|_{L}\times\nu)|_{Q_{3}(\mathbf{0})},$ for some $2$-ADR measure $\nu$ on $\mathbb{R}$, which yields that $\alpha_{k,\mu}(B_{1}(\mathbf{0}))=0$, a contradiction. Now, depending on whether or not there is a further orthogonal vector $e_{2}$, either we can apply Lemma 5.12 to get that $\alpha_{k+1,\mu}(\mathbf{0},M_{1})<\lambda$, or Lemma 5.12 which yields that either $\alpha_{k+1,\mu}(\mathbf{0},M_{1})<\lambda$ or $\alpha_{k+2,\mu}(\mathbf{0},M)<\lambda$ for every $M\geq M_{2}$.
Next suppose $\dim L=k+1$. If $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{5M_{1}}(\mathbf{0})=L$, then Lemma 5.9 ensures that $\mu|_{Q_{M_{1}}(\mathbf{0})}=(\mathcal{H}^{k+1}|_{L}\times\nu)|_{Q_{M_{1}}(\mathbf{0})}$ for some $1$-ADR measure $\nu$ on $\mathbb{R}$, which yields that $\alpha_{k+1,\mu}(\mathbf{0},M_{1})=0$. Otherwise, $\operatorname{supp}(\mu)_{0}\cap\widetilde{Q}_{5M_{1}}(\mathbf{0})\neq L$. There exists $e_{1}\in\operatorname{supp}(\mu)_{0}$ orthogonal to $L$ which satisfies the conclusions of Lemma 5.11. Notice that, because part (2)(b) of Lemma 5.11 holds, we may assume that $|e_{1}|<5M_{1}$. In this case we apply Lemma 5.12 to $\widetilde{\mu}=\frac{\mu(\,\cdot\,\,5M_{1})}{(5M_{1})^{k+2}}$ and $\widehat{e}_{1}=(5M_{1})^{-1}e_{1}$ to conclude $\alpha_{k+1,\widetilde{\mu}}(\mathbf{0},M)<\lambda$ for every $M\geq M_{1}$. This immediately implies $\alpha_{k+2,\mu}(\mathbf{0},5MM_{1})<\lambda$ for every $M\geq M_{1}$.
Finally, suppose $\dim L=k+2$. Then we must have $\operatorname{supp}(\mu)_{0}=L$, and so Lemma 5.9 ensures that $\mu=\mathcal{H}^{k+1}_{|L}\times\nu$ and hence that $\alpha_{k+2,\mu}(\mathbf{0},r)=0$ for any $r>0$.
Selecting $\widehat{M_{2}}\geq\max(M_{2},5M_{1}^{2})$ now yields that, in every case, either
$\alpha_{k+1,\mu}(\mathbf{0},M_{1})<\lambda\text{ or }\alpha_{k+2,\mu}(\mathbf{0},\widehat{M}_{2})<\lambda,$
and the proposition is proved. ∎
@@ 5.2. The Heisenberg group case
$\Sigma_{\mathbf{X}}(\mathbf{Y})=(2X-Y,t_{Y}+Y_{1}X_{2}-X_{1}Y_{2})$
@@
$=\mu(\mathbf{X}\cdot B_{r}(Y-X,t_{Y}-t_{X}+\frac{1}{2}(Y_{1}X_{2}-X_{1}Y_{2}))$
@@
Orponen [51] called a Borel measure $\mu$ on $\mathbb{H}$ symmetric measure if, for each pair $\mathbf{X},\mathbf{Y}\in\operatorname{supp}(\mu)$, $\Sigma_{\mathbf{X}}(\mathbf{Y})\in\operatorname{supp}(\mu)$, and proved the following result:
@@
[51] The support of a $3$-regular, symmetric measure on $\mu$ on $\mathbb{H}$ is contained in a vertical plane.
@@
Suppose $\mu$ is a $3$-regular, vampiric measure on $\mathbb{H}$ whose support contains $(0,0,0)$. Then $\mu=\mathcal{H}^{1}|_{L}\times\nu$ for some line $L\subset\mathbb{R}^{2}$ and a $2$-ADR measure $\nu$ on $\mathbb{R}.$ Conversely, any measure of the form $\mu=\mathcal{H}^{1}|_{L}\times\nu$ is a vampiric measure on $\mathbb{H}$.
@@
Orponen’s theorem (Corollary 5.16 ensures that, for some vertical hyperplane $L\times\mathbb{R}\subset\mathbb{H}$, $\operatorname{supp}(\mu)\subset L\times\mathbb{R}$. If $\mathbf{X},\mathbf{Y}\in L$, then $\mathbf{X}\cdot\mathbf{Y}=(X+Y,t_{X}+t_{Y})$, and so the Heisenberg group action becomes the Parabolic group action on $L$. Therefore Lemma 5.11 ensures that $(\operatorname{supp}(\mu))_{0}=L$. Appealing now to Lemma 5.9 completes the proof (the converse direction is a trivial computation.) ∎
@@ 6. Packing bad cubes
Our main proposition is the following. Recall the quantities $M_{1}(\lambda)$ and $M_{2}(\lambda)$ depending on $\lambda$ in Proposition 5.2.
@@
(6.1) $\alpha_{k,\mu}(\mathbf{Z}_{E},\ell(E))\geq\varkappa,\;\alpha_{k+1,\mu}(\mathbf{Z}_{E},M_{1}(\lambda)\ell(E))\geq\lambda,\,\text{ and }\,\alpha_{k+2,\mu}(\mathbf{Z}_{E},M_{2}(\lambda)\ell(E))\geq\lambda,$
@@
$\max_{\varphi\in\mathcal{F}}\Theta_{\mu,\varphi,A}(E)\geq\Delta\mu(E).$
@@
Suppose that $\mu$ is $(k+2)$-ADR, and there exists $\varepsilon$ such that $\varepsilon$-$\bf HSDC$} cubes are Carleson. Then there exists $\lambda>0$ such that the collection of cubes $E\in\mathbb{D}(\mu)$ such that
$\alpha_{k+1,\mu}(\mathbf{Z}_{E},M_{1}\ell(E))<\lambda\text{ or }\alpha_{k+2,\mu}(\mathbf{Z}_{E},M_{2}\ell(E))<\lambda$
@@
Suppose that $\mu$ is a $(k+2)$-ADR measure for which is good for all spatially asymmetric Littlewood-Paley kernels, and the set of $\varepsilon$-$\bf HSDC$} cubes are Carleson for some $\varepsilon>0$. Fix $\varkappa$ to be equal to be the value of $\lambda$ in Theorem 4.3. Select $\lambda$ to be as in Lemma 6.2. First, using Lemma 2.5, we find that, the collection cubes $E\in\mathbb{D}(\mu)$ which satisfy (6.1) are Carleson. But now, since the $\varepsilon$-HSDC cubes are Carleson, Proposition 6.1 and Lemma 6.2 ensure that the collection of cubes $E\in\mathbb{D}(\mu)$ such that $\alpha_{k,\mu}(\mathbf{Z}_{E},\ell(E))>\varkappa$ is Carleson. But now Theorem 4.3 ensures that $\mu$ is uniformly recitifiable.∎
@@
Fix $\varkappa>0$ and $\lambda>0$. Let $\{\varphi_{j}\}_{j}$ be a countable, dense subset of $\operatorname{Lip}_{0}^{\operatorname{odd}}(\mathbb{R}^{n+1})$ with $\|\varphi\|_{\operatorname{Lip}}\leq 1$, such that for each $j$, $\varphi_{j}\in C^{\infty}(\mathbb{R}^{n+1})$ and $\operatorname{supp}(\varphi_{j})\subseteq B_{j}(\mathbf{0})$ for each $j$.
@@
[…]
@@
$\max_{j\in\{1,2,...,\ell\}}\Theta^{\mu_{\ell}}_{\varphi_{j},\ell}(E_{\ell})\leq\frac{1}{\ell}$
@@
$\alpha_{\mu,k}(\mathbf{Z}_{E},\ell(E))>\lambda,$